# Encoding of lexical tone in self-supervised models of spoken language

### v1

@@ 1 Introduction
Explaining the inner workings of self-supervised models of written and spoken language has been the focus of much recent work. Transformer-based (Vaswani et al., 2017) written language models have been shown to encode many types of linguistic information (Conneau et al., 2018; Hewitt and Manning, 2019). The analysis of self-supervised Spoken Language Models (SLMs) is also gaining traction: architectures such as wav2vec2 (Baevski et al., 2020) and HuBERT (Hsu et al., 2021) have been shown to encode linguistic information at the phonetic, phonological, syntactic and semantic levels of human speech without labeled data (Abdullah et al., 2021; Ma et al., 2021; de Seyssel et al., 2022; Bartelds et al., 2022; Martin et al., 2023; Shen et al., 2023).
The majority of research on representations of phonetic and phonological information in SLMs focuses on the segmental level. Segmental refers to units of speech that do not spread but remain localized. Phonemes (e.g. vowels and consonants) are the smallest abstract units of sound that help to distinguish one unit from another (e.g. pat vs bat). Suprasegmental, in contrast, refers to features that are not necessarily limited to single units, but can spread across multiple phonemes or phrases. Examples include tone, stress patterns, and intonation, which can all entail syllable and phrase level changes Singh and Fu (2016). The representation of suprasegmental information in SLMs is important to study, as it is one of the main distinguishing features of speech compared to text: spoken utterances use suprasegmental cues to convey information that is sometimes not explicitly marked in a corresponding written sentence [^1] Mandarin does not mark tone in transcription, but Vietnamese does.. As a first step, in this work, we focus on lexical tone as a highly constrained, relatively well-understood example of a suprasegmental feature.
@@ 2 Tones
The Tone Bearing Unit (TBU) is the segment containing tone; this is typically, although not always, found on the syllable level. In the case of the primary language investigated in this paper, Mandarin, syllables are morpheme based (as in, Mandarin syllables are morphemes). Thus in Mandarin every morpheme contains obligatory tone, and is in of itself the TBU. (Jun and Kubozono, 2020).
@@
We compare SLMs trained on non-tonal language as well as three fully lexical tonal languages: Mandarin, Cantonese and Vietnamese. The models are tested primarily on Mandarin data. Mandarin demonstrates full tonality, with tone found on each morpheme (Hyman, 2018), and has been widely studied for tone perception as well as acquisition. Secondarily, we also test on data from another lexical tone language, Vietnamese, to assess if our results generalize.
@@ 4 Methodology
To examine the encoding of tone, we train linear probing classifiers on the hidden state activations extracted from the aforementioned models for every morpheme in our testing datasets.  [^2] We release the codebase for our experiments https://github.com/techsword/tone-encoding-in-speech-model
@@ 5.1 Tone encoding across languages
We repeat the tone classification experiment for Vietnamese tones. Results in Figure 3 show the Cantonese model performs slightly better than the English model, especially towards the later layers; the Mandarin model, however, patterns similar to the English model. This is could be due to the fact that Mandarin has fewer tonal contrasts than Vietnamese and Cantonese (cf. Section 2), or, more likely, that Vietnamese uses different acoustic cues such as phonation type and voice quality to in tonal perception than f0 contours or height in Mandarin (Brunelle, 2009).
@@ 5.3.2 Tone and consonant contrasts
Non-native speakers can have difficulty distinguishing between T2-T3 and T1-T4 tone pairs in Mandarin (Hao, 2012). While native adult listeners have shown near perfect tone identification accuracy (So and Best, 2010; Tsukada and Kondo, 2019), there is literature documenting native Mandarin-learning toddlers and also some adults having more difficulty in distinguishing tone pair T2-3 compared to other combinations of Mandarin lexical tone. Huang and Johnson (2011) showed that tone pair T2-3 is the most confusable for both native speakers of American English and Mandarin while T1-4 also being confusable for native Mandarin speakers. We investigate this pattern in pre-trained SLMs via a dedicated probing experiment, using the final (85,000 steps) checkpoint of the pre-trained models in Section 5.3.1. As can be seen in Figure 7, tone pairs T1-4 and T2-3 show the largest differences in the best classification accuracy between the Mandarin and English models, which roughly matches human perceptual pattern.
@@ 6 Conclusion
We analyze the tone encoding capabilities of spoken language models trained on three tonal and two non-tonal languages, using classifier probes with data from two tonal languages Mandarin and Vietnamese. We find that SLMs trained on either tonal or non-tonal languages encode tonal information in Mandarin and Vietnamese to a significant degree.

### v2

@@ 1 Introduction
Explaining the inner workings of self-supervised models of written and spoken language has been the focus of much recent work. Transformer-based (Vaswani et al., 2017) written language models have been shown to encode many types of linguistic information (Conneau et al., 2018; Hewitt and Manning, 2019). The analysis of self-supervised Spoken Language Models (SLMs) is also gaining traction: architectures such as wav2vec2 (Baevski et al., 2020) and HuBERT (Hsu et al., 2021) have been shown to encode linguistic information at the phonetic, phonological, syntactic and semantic levels of human speech without labeled data (Abdullah et al., 2021; Ma et al., 2021; de Seyssel et al., 2022; Bartelds et al., 2022; Martin et al., 2023; Shen et al., 2023; Pasad et al., 2024).
The majority of research on representations of phonetic and phonological information in SLMs focuses on the segmental level. Segmental refers to units of speech that do not spread but remain localized. Phonemes (e.g. vowels and consonants) are the smallest abstract units of sound that help to distinguish one unit from another (e.g. pat vs bat). Suprasegmental, in contrast, refers to features that are not necessarily limited to single units, but can spread across multiple phonemes or phrases. Examples include tone, stress patterns, and intonation, which can all entail syllable and phrase level changes Singh and Fu (2016). The representation of suprasegmental information in SLMs is important to study, as it is one of the main distinguishing features of speech compared to text: spoken utterances use suprasegmental cues to convey information that is sometimes not explicitly marked in a corresponding written sentence. [^1] Mandarin does not explicitly mark tone in writing, but Vietnamese does. As a first step, in this work, we focus on lexical tone as a highly constrained, relatively well-understood example of a suprasegmental feature.
@@ 2 Tones
The Tone Bearing Unit (TBU) is the segment containing tone; this is typically, although not always, found on the syllable level. In the case of the primary language investigated in this paper, Mandarin, syllables are morpheme based (as in, Mandarin syllables are morphemes). Thus in Mandarin every morpheme contains obligatory tone, and is in of itself the TBU (Jun and Kubozono, 2020).
@@
We compare SLMs trained on non-tonal languages as well as three fully lexical tonal languages: Mandarin, Cantonese and Vietnamese. The models are tested primarily on Mandarin data. Mandarin demonstrates full tonality, with tone found on each morpheme (Hyman, 2018), and has been widely studied for tone perception as well as acquisition. Secondarily, we also test on data from another lexical tone language, Vietnamese, to assess if our results generalize.
@@ 4 Methodology
To examine the encoding of tone, we train linear probing classifiers on the hidden state activations extracted from the aforementioned models for every morpheme in our testing datasets.  [^2] We release the codebase for our experiments at https://github.com/techsword/tone-encoding-in-speech-model
@@ 5.1 Tone encoding across languages
We repeat the tone classification experiment for Vietnamese tones. Results in Figure 3 show the Cantonese model performs slightly better than the English model, especially towards the later layers; the Mandarin model, however, patterns similar to the English model. This is could be due to the fact that Mandarin has fewer tonal contrasts than Vietnamese and Cantonese (cf. Section 2), or, more likely, that Vietnamese uses different acoustic cues such as phonation type and voice quality in tonal perception than f0 contours or height in Mandarin (Brunelle, 2009).
@@ 5.3.2 Tone and consonant contrasts
Non-native speakers can have difficulty distinguishing between T2-T3 and T1-T4 tone pairs in Mandarin (Hao, 2012). While native adult listeners have shown near perfect tone identification accuracy (So and Best, 2010; Tsukada and Kondo, 2019), there is literature documenting native Mandarin-learning toddlers and also some adults having more difficulty in distinguishing the tone pair T2-3 compared to other combinations of Mandarin lexical tone. Huang and Johnson (2011) showed that the tone pair T2-3 is the most confusable for both native speakers of American English and Mandarin while T1-4 also being confusable for native Mandarin speakers. We investigate this pattern in pre-trained SLMs via a dedicated probing experiment, using the final (85,000 steps) checkpoint of the pre-trained models in Section 5.3.1. As can be seen in Figure 7, tone pairs T1-4 and T2-3 show the largest differences in the best classification accuracy between the Mandarin and English models, which roughly matches human perceptual pattern.
@@ 6 Conclusion
We analyze the tone encoding capabilities of spoken language models trained on three tonal and two non-tonal languages, using classifier probes with data from two tonal languages: Mandarin and Vietnamese. We find that SLMs trained on either tonal or non-tonal languages encode tonal information in Mandarin and Vietnamese to a significant degree.